The weight of a boat varies directly with the length of the boat. If a 12 foot boat weighs 86 lbs. then what is the weight of a 14 foot boat? (Find the weight to the nearest tenth of a pound.)

Answer by checkley77(12844)

12/86=14/x
12x=86*14
12x=1,204
x=1,204/12
x=100 pounds.